# ChartAnchor: Chart Grounding with Structural-Semantic Fidelity and Data Recovery

### v1

@@
Abstract.
Recent advances in multimodal large language models (MLLMs) highlight the need for benchmarks that rigorously evaluate structured chart comprehension. Chart grounding refers to the bidirectional alignment between a chart’s visual appearance and the structured semantics. This task requires models to produce a symbolic specification that faithfully captures the chart’s visual and structural intent, while also recovering the underlying tabular data with precise values and relationships. Chart grounding directly reflects a model’s capabilities in numerical reasoning, multimodal alignment, and structural reconstruction, and has several important applications in real-world scenarios. Existing benchmarks, constrained by narrow chart diversity, isolated tasks, and incomplete evaluation frameworks, fail to holistically assess grounding. To address this, we propose ChartAnchor, a comprehensive benchmark of 8k+ chart-table-code triples spanning 30 chart types drawn from diverse real-world and augmented sources. ChartAnchor introduces two complementary tasks: chart-to-code generation (synthesizing executable code to replicate charts) and controlled chart-to-table reconstruction (extracting exact data with predefined headers), enabling cross-validation of visual and numerical fidelity. A multi-level evaluation framework integrates semantic validation, stylistic analysis, and perceptual metrics to assess both structural and content-level correctness. Extensive experiments on MLLMs reveal critical limitations in numerical precision and code synthesis, emphasizing the need for structured reasoning beyond surface-level perception. By unifying symbolic and data-driven grounding, ChartAnchor establishes a rigorous foundation for chart grounding, offering meaningful insights for advancing MLLMs in scientific, financial, and industrial domains. Our code and data are available at https://github.com/immortal5655/ChartAnchor.
Chart Grounding; Chart Understanding; MLLMs; Multimodal Reasoning.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM SIGKDD Conference on Knowledge Discovery and Data Mining; June 03–05, 2026; Jeju Island, South Koreaisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
While recent chart-related benchmarks such as visual QA (e.g., ChartQA(masry2022chartqa), PlotQA(methani2020plotqa)) and the summarization (e.g., Chart-to-Text(kantharaj2022chart), ChartSumm (rahman2023chartsumm)) have advanced chart understanding, they evaluate only unstructured outputs—offering limited insight into a model’s ability to recover symbolic or structured content. Tasks like chart-to-table and chart-to-code begin to address structured aspects, but each only tells part of the story. Chart-to-table benchmarks assess data recovery but ignore style and structure. Conversely, chart-to-code (e.g. ChartMimic(yang2024chartmimic), Plot2Code(wu2024plot2code)) benchmarks are predominantly focus on visual appearance, including layout, styling, and chart-surface text. While they may assess textual accuracy on the chart surface, these benchmarks still judge success primarily through visual similarity and do not enforce consistency with the underlying structured data, which poses serious risks in data-sensitive scenarios.
@@
2. Related Work
MLLM. Building on the success of Large Language Models (LLMs) (touvron2023llama; bai2023qwen), Multimodal Large Language Models (MLLMs) extend their capabilities to jointly process visual, video, and textual data, showing state-of-the-art performance across diverse scenarios. The current landscape can be divided into two branchs: 1) Open-source models—such as mPLUG-Owl2 (ye2024mplug), Qwen2.5-VL(wang2024qwen2), CogVLM (hong2024cogvlm2), and the more recent InternVL3 (zhu2025internvl3) employ efficient architectures through frozen vision encoders, lightweight adapters, or mixture-of-experts (MoE) designs. These models target specialized capabilities such as long-context processing, document analysis, and multilingual support, facilitating deployment in resource-limited scenarios. 2) Closed-source models (e.g., GPT-4V (yang2023dawn), GPT-4o (openai2024) Gemini-1.5-Pro (team2024gemini), Claude-3 (anthropic2024haiku)), leverage vast private datasets and extreme computational budgets to fuse text, image, audio, and even video streams within a single architecture. While their internal parameters are not publicly available, these systems establish widely adopted standards for academic research and industry applications, advancing evaluation methodologies and safety studies.
Chart Benchmark. As chart comprehension emerges as a key challenge for multimodal reasoning in MLLMs, a variety of benchmarks have been developed to assess different capabilities. Chart-to-Text (kantharaj2022chart) and ChartSumm (rahman2023chartsumm) evaluate summarization via structured textual descriptions. For visual question answering, ChartQA (masry2022chartqa), FigureQA (kahou2017figureqa), DVQA (kafle2018dvqa) and MultiChartQA (zhu2024multichartqa) provide multi-scale challenges spanning element localization and trend extrapolation. Benchmarks like ChartMimic (yang2024chartmimic), Plot2Code (wu2024plot2code), and Design2Code (si2024design2code) assess models’ abilities to generate accurate and executable code from chart images, highlighting the challenges in aligning visual understanding with code generation. While recent benchmarks such as ChartBench (xu2023chartbench) and ChartX (xia2024chartx) move beyond isolated tasks by incorporating diverse reasoning and perception challenges, they do not explicitly unify visual understanding with structured semantic outputs. Moreover, existing datasets are limited by their diversity in chart types, fixed rendering libraries, and metric calculations.
In contrast to existing benchmarks, we introduce ChartAnchor, a new benchmark designed to advance the field by formalizing the chart grounding task. Chart grounding is a unified task that jointly maps visual charts to both executable code and tabular data, effectively bridging chart-to-code generation and controlled table extraction within a single, interpretable framework. Despite its practical relevance, this task has been largely overlooked by existing benchmarks. ChartAnchor includes 30 diverse chart types, over 8,000 test instances, and rich image–table–code tuples rendered from multiple charting libraries, enabling more robust and generalizable evaluation of MLLMs.
Figure 1. Chart Type Distribution in ChartAnchor.
@@ 2. Related Work
Figure 2. Illustration of our chart-to-code generation task and controlled chart-to-table reconstruction task.
3. The ChartAnchor Benchmark
To support robust evaluation of chart grounding, this section introduces two core tasks: chart-to-code generation and controlled chart-to-table reconstruction. We construct a diverse corpus of chart–table–code triplets by leveraging both existing datasets and real-world sources. The overall distribution of chart types in ChartAncor is shown in Fig. 2. We then present our parametric code generation and augmentation pipeline, followed by a filtering process that combines automated validation and expert review to ensure high-quality, semantically faithful data.
3.1. Task Definitions
As depicted in Fig. 2, to support robust evaluation of chart grounding, we define two core tasks in our benchmark: chart-to-code generation and controlled chart-to-table reconstruction. Each task is grounded in the chart–table–code triplets we collect and reflects distinct practical scenarios relevant to real-world chart understanding and visual reasoning.
Chart-to-Code Generation. This task challenges a model to generate executable Python plotting code from a chart image. Given an input image $I$, the model must produce a valid script that accurately reconstructs the chart’s visual and structural elements, including the chart type, underlying data, axes, titles, legends, color scheme, and layout. The generated code should be both syntactically correct and semantically faithful, yielding a chart that closely matches the input. This task assesses the model’s capabilities in visual abstraction, symbolic reasoning, and code synthesis.
Controlled Chart-to-Table Reconstruction. This task focuses on recovering the underlying tabular data from a chart image under constrained header supervision. Given an input chart image $I$ and a predefined set of column headers $\mathcal{H}=\{h_{1},h_{2},\dots,h_{n}\}$, the model is required to generate a table $\mathcal{T}=[r_{1},r_{2},\dots,r_{m}]$, where each row $r_{j}$ assigns numerical values to the corresponding headers. This task evaluates the model’s ability to perform fine-grained numerical reasoning and precise visual interpretation. The inclusion of headers removes label ambiguity, enabling a more targeted and reliable evaluation of data fidelity.
3.2. Data Collection
Figure 3. Illustration of the Data Collection Pipeline.
@@ 3.2. Data Collection
Table 1. Dataset statistics.
In this section, we briefly present our data collection pipeline for building a high-quality chart–table–code corpus. ChartAnchor comprises 8,068 validated samples, with 1,535(19%) from augmented existing datasets and 6,533(81%) curated from real-world sources. Detailed procedures are described in Appendix A.
3.2.1. Data Source
  • Existing datasets. We collect chart–table pairs from four well-established datasets: PlotQA(methani2020plotqa), DVQA(kafle2018dvqa), FigureQA(kahou2017figureqa) and Vistext(tang2023vistext). Since these datasets do not provide the corresponding Python code, we construct scripts to regenerate the chart images using the provided tables and chart metadata. To enhance diversity and prevent data leakage, we further apply systematic augmentations to various visual and semantic attributes during code generation.
3.2.2. Code Generation and Augmentation for Existing Data-
sets.
3.2.3. Filtering.
Our filtering process consists of two stages:
Automated Filtering. We apply a four-stage filtering process to ensure data quality and consistency:
  • (a) Completeness Check: Remove all incomplete data (table, chart, code) triples.
  • (b) Structural Filtering: Exclude geospatial maps, rasterized tables, and overly decorated charts that lack recoverable structure.
  • (c) Deduplication: We compute a compact feature signature for each sample based on data characteristics, including the number of entries, value counts, mean, standard deviation, min/max values, and the number of unique values. These features are encoded as string signatures for fast comparison, and duplicates are filtered by exact signature matching.
  • (d) Executability Check: We execute all Python scripts and discard any triples with runtime errors or failed rendering.
Human Filtering. Each chart–table–code triple was independently reviewed by experts experienced in deep learning and with hands-on experience in Python programming, data visualization, and chart scripting. Samples were grouped by chart type and filtered within each group to ensure consistency across chart families. Using a standardized rubric, reviewers evaluated each triple along three dimensions: semantic accuracy, visual clarity, and stylistic diversity, and assigned per-triple scores accordingly. Each sample was reviewed by two experts under a blind, rotating setup. In cases of major disagreement, a third expert served as an adjudicator to ensure annotation reliability and consensus.
(a) Code Length Distribution.
(b) Image Entropy Distribution.
(c) Brightness vs. Contrast.
(d) Visual vs. Color Complexity.
Figure 4. Chart-level visual statistics.
@@ 3.2.2. Code Generation and Augmentation for Existing Data-
Table 2. Chart counts across types.
3.3. Dataset Analysis
@@
Dataset Types
Test
Set
Input
Format
Output
Format
Multi-task
Eval
Visual
Open
Source
Plot
Library
Data
Eval
Full
Ground
Chart Benchmarks
ChartQA(masry2022chartqa) 3 10k I+NL NL - -
PlotQA(methani2020plotqa) 3 34k I+NL NL - -
Chart-to-text(kantharaj2022chart) 6 44k I+NL NL - -
OpenCQA(kantharaj2022opencqa) 5 1.2k I+NL NL - -
ChartSumm(rahman2023chartsumm) 3 84k I+NL NL - -
Charxiv(wang2024charxiv) 18 1.32k I+NL NL - -
MMC(liu2023mmc) 6 2k I+NL NL - -
ChartX(xia2024chartx) 18 6k I+NL NL - -
ChartBench(xu2023chartbench) 9 2.1k I+NL NL - -
Code Generation Benchmarks
HumanEval(chen2021evaluating) - 164 Code Code - - - -
MBPP(austin2021program) - 500 NL+Code Code - - - -
MMCode(li2024mmcode) - 263 I+NL Code - - - -
MatPLotBench(yang2024matplotagent) 13 100 NL Code Mixture
PLot2Code(wu2024plot2code) 15 132 I+NL Code Singal
ChartMimic(yang2024chartmimic) 22 4.8k I+NL Code Singal
Design2Code(si2024design2code) - 484 I+NL Code Singal
ChartAnchor 30 8.1k I+NL Code+NL Mixture
Table 3. Comparison of benchmarks for chart understanding and code generation. “I” = image, “NL” = natural language. “Visual” assesses unannotated chart reasoning. “Plot Library” denotes the use of multiple plotting libraries in the chart-to-code task, while “Data Eval” indicates whether data fidelity is evaluated within the chart-to-code task.
As shown in Tab. 2, ChartAnchor encompasses 30 visualization families. Beyond canonical 2D Cartesian plots (e.g., bar, line, scatter), it cover a wide variety of chart types: (i) advanced 3D charts such as scatter3d, surface, mesh3d, and line3d; (ii) hierarchical/flow diagrams including sunburst, sankey, treemap, and funnelarea; (iii) polar and ternary variants (e.g., scatterpolar, barpolar); and (iv) finance–specific charts such as candlestick and ohlc. The overall distribution reflects realistic usage patterns, with frequent types (bar, line, pie) having 600–700 samples, mid-frequency ones 200–400, and fewer complex charts. By combining curated benchmarks with real-world user code, ChartAnchor exposes models to a diverse set of plotting libraries (e.g., matplotlib, plotly), varied scripting idioms, and domain-specific styling practices, thereby offering a robust and practical benchmark for evaluating the chart grounding capabilities of MLLMs. More analysis are in Appendix B. We also provide representative chart examples from each category, highlighting visual diversity and structural complexity, are illustrated in Appendix G.
3.4. Comparition with other benchmarks
Table 3 presents a comprehensive comparison between our proposed benchmark and a wide range of existing datasets related to chart understanding and code generation. We evaluate them along key dimensions, including input/output modalities, task diversity, support for visual reasoning, openness, plotting library variety, data fidelity assessment, and the completeness of symbolic supervision.
Earlier chart benchmarks such as ChartQA, PlotQA, and Chart-to-Text primarily focus on natural language tasks like question answering and summarization, without structured protocols to evaluate bidirectional alignment between chart visuals and their underlying data or specifications. More recent multi-task datasets such as ChartArxiv and ChartBench broaden the evaluation scope, yet still omit code-based grounding, limiting their ability to assess executable understanding of visual semantics.
Code generation benchmarks like HumanEval and MBPP are designed for text-to-code synthesis and lack any visual input, making them unsuitable for evaluating multimodal alignment. Multimodal code-focused datasets like Plot2Code and ChartMimic introduce image-to-code generation but do not explicitly assess numerical fidelity or data reconstruction accuracy.
@@ 4. Evaluation Metrics
  • 1) Textual Components Match. We extract textual components—titles, axis labels, legends, annotations—and compare them against reference charts to verify semantic and positional consistency.
    Text Color Type Layout
    Legend Title
    Axis
    Label
    Annos Avg.
    Proprietary Multimodal Large Language Models
    GPT4o 91.88 63.53 72.83 67.20 76.36 69.98 34.64 70.06 80.73 35.85 86.88 63.02
    Claude-3-7-Sonnet 78.60 54.97 69.11 67.19 70.57 65.46 40.79 59.89 76.53 30.30 94.11 61.18
    Open-Source Multimodal Large Language Models
    InternVL3-2B 28.97 18.75 17.41 17.63 20.88 18.67 8.98 16.01 21.55 13.56 26.95 17.62
    Qwen2.5-VL-3B-Instruct 48.07 28.24 31.80 26.86 39.44 31.58 12.29 27.01 40.19 16.71 43.98 28.63
    Gemma-3-4B-it 66.20 42.29 38.72 36.41 57.50 43.73 18.74 33.53 60.46 24.21 60.98 40.28
    DeepSeek-VL-7B 32.05 17.59 20.41 17.62 23.19 19.70 9.56 17.37 24.83 7.74 31.34 18.42
    LLaVA-v1.6-Mistral-7B 15.32 8.74 8.37 4.96 4.91 6.75 6.03 7.79 5.05 4.76 13.94 7.39
    Qwen2.5-VL-7B-Instruct 67.59 41.75 41.46 43.60 61.90 47.18 18.45 41.83 64.30 27.85 62.48 43.68
    MiniCPM-V-2.6-8B 26.87 16.76 16.88 14.21 16.18 16.01 7.91 13.77 16.92 12.95 24.87 15.40
    InternVL3-9B 69.19 44.75 47.48 45.06 59.57 49.20 22.06 40.72 66.20 29.20 64.34 45.29
    GLM-4V-9B 46.04 26.55 27.86 24.55 43.78 30.68 12.02 23.32 45.69 15.51 42.60 28.30
    CogVLM2-Llama3-Chat-19B 7.82 7.12 6.03 4.79 1.25 4.80 3.77 5.61 1.21 3.52 7.17 4.35
    InternVL3-14B 80.60 52.80 55.69 54.90 75.12 59.63 26.13 53.08 80.63 30.59 75.49 54.26
    Qwen2.5-VL-32B-Instruct 67.59 42.71 48.69 48.91 73.55 53.47 23.56 48.29 82.08 21.14 64.04 48.76
    Table 4. Comprehensive performance comparison of proprietary and open-source MLLMs on the Chart-to-Code generation task. For data fidelity, we report IoU score under a slight tolerance stage. As our collected matplotlib images do not include annotations by design, they are excluded from the annotation metric evaluation.
  • 2) Color Fidelity. We quantify perceptual color differences using the CIEDE2000 (sharma2005ciede2000) metric in the CIE Lab color space, which aligns with human vision sensitivity. This evaluation covers both static elements (background, axes) and dynamic color mappings (data series, gradients). For multi-color comparisons (e.g., categorical series), we apply the Hungarian algorithm to optimally pair generated and reference colors, minimizing the total perceptual deviation ($\Delta E_{00}$) across matched pairs. across matched pairs. This ensures semantic alignment of color associations while preserving numerical fidelity.
@@
To evaluate whether models can faithfully recover the underlying chart data, we introduce a data-level fidelity metric, applicable to both chart-to-code and controlled chart-to-table tasks. To the best of our knowledge, this is the first work to introduce structured data fidelity evaluation in the chart-to-code setting, moving beyond functional or visual correctness to explicitly verify the semantic integrity of generated chart data. In the chart-to-code task, we first parse the generated figure objects and dispatch them to type-specific extractors $\text{Extract}_{T}$, where each $T\in\mathcal{T}$ corresponds to a supported chart type (e.g., line, bar, candlestick). Each extractor retrieves the relevant data fields—such as $(x,y)$ for line or $(x,\text{low},\text{high},\text{open},\text{close})$ for candlestick charts—and formats each data point as a normalized tuple $\tau_{i}=(n,x_{i},y_{i},\ldots)$, where $n$ is the trace name and the remaining elements are type-specific field values. The final structured form is a set $\mathcal{L}_{T}=\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$. In the controlled chart-to-table task, each table row is similarly treated as a tuple $\tau_{i}=(t_{1},t_{2},\ldots,t_{m})$, resulting in a comparable structured set $\mathcal{L}_{T}$, allowing both tasks to be evaluated under a unified tuple-based framework.

### v3

@@
Abstract
Recent advances in multimodal large language models (MLLMs) highlight the need for benchmarks that rigorously evaluate structured chart comprehension. Chart grounding refers to the bidirectional alignment between a chart’s visual appearance and the structured semantics. This task requires models to produce a symbolic specification that faithfully captures the chart’s visual and structural intent, while also recovering the underlying tabular data with precise values and relationships. Chart grounding directly reflects a model’s capabilities in numerical reasoning, multimodal alignment, and structural reconstruction, and has several important applications in real-world scenarios. Existing benchmarks, constrained by narrow chart diversity, isolated tasks, and incomplete evaluation frameworks, fail to holistically assess grounding. To address this, we propose ChartAnchor, a comprehensive benchmark of 8k+ chart-table-code triples spanning 30 chart types drawn from diverse real-world and augmented sources. ChartAnchor introduces two complementary tasks: chart-to-code generation and controlled chart-to-table reconstruction, enabling cross-validation of visual and numerical fidelity. A multi-level evaluation framework integrates semantic validation, stylistic analysis, and perceptual metrics to assess both structural and content-level correctness. Extensive experiments on MLLMs reveal critical limitations in numerical precision and code synthesis, emphasizing the need for structured reasoning beyond surface-level perception. By unifying symbolic and data-driven grounding, ChartAnchor establishes a rigorous foundation for chart grounding, offering meaningful insights for advancing MLLMs in scientific, financial, and industrial domains.
ChartAnchor: Chart Grounding with Structural-Semantic Fidelity and Data Recovery
Xinhang Li, Jingbo Zhou, Pengfei Luo, Yixiong Xiao, Tong Xu Baidu Research, USTC {lixinhang, zhoujingbo, luopengfei, xiaoyixiong}@baidu.com, tongxu@ustc.edu.cn
1 Introduction
@@
While recent chart-related benchmarks such as visual QA (e.g., ChartQA(Masry et al., 2022), PlotQA(Methani et al., 2020)) and the summarization (e.g., Chart-to-Text(Kantharaj et al., 2022b), ChartSumm (Rahman et al., 2023)) have advanced chart understanding, they evaluate only unstructured outputs—offering limited insight into a model’s ability to recover symbolic or structured content. Tasks like chart-to-table and chart-to-code begin to address structured aspects, but each only tells part of the story. Chart-to-table benchmarks assess data recovery but ignore style and structure. Conversely, chart-to-code (e.g. ChartMimic(Yang et al., 2024b), Plot2Code(Wu et al., 2024a)) benchmarks are predominantly focus on visual appearance, including layout, styling, and chart-surface text. While they may assess textual accuracy on the chart surface, these benchmarks still judge success primarily through visual similarity and do not enforce consistency with the underlying structured data, which poses serious risks in data-sensitive scenarios.
@@
2 Related Work
Figure 1: Chart Type Distribution in ChartAnchor.
@@ 2 Related Work
Figure 2: Illustration of our chart-to-code generation task and controlled chart-to-table reconstruction task.
MLLMs.
Proprietary pioneers, notably Claude-Sonnet-4 (Anthropic, 2025b), GPT-5 (OpenAI, 2025), and Gemini-2.5-Pro (Comanici et al., 2025), have unlocked potential for addressing complex realistic applications. Concurrently, the open-source community has released diverse high-quality models. This includes prominent series like QwenVL (Yang et al., 2025) and InternVL (Zhu et al., 2025), alongside models such as MiMo-VL (Xiaomi and Team, 2025), GLM-4V (GLM et al., 2024a), and Deepseek-VL (Lu et al., 2024). Together, they significantly advance the field of visual-textual understanding.
Chart Benchmark. The growing importance of chart comprehension in multimodal reasoning has driven the development of numerous benchmarks. Existing datasets target distinct capabilities, ranging from summarization (Kantharaj et al., 2022b; Rahman et al., 2023) and VQA (Masry et al., 2022; Methani et al., 2020; Zhu et al., 2024) to code generation (Yang et al., 2024b; Wu et al., 2024a; Si et al., 2024). Although ChartBench (Xu et al., 2023) and ChartX (Xia et al., 2024) move beyond isolated tasks by incorporating diverse reasoning challenges, they fail to unify visual understanding with structured semantic outputs. Moreover, existing datasets are limited by their diversity in chart types, fixed rendering libraries, and rigid metrics.
In contrast to existing benchmarks, we introduce ChartAnchor, a benchmark that formalizes chart grounding—a unified task mapping visual inputs to both executable code and tabular data. ChartAnchor bridges bridging chart-to-code generation and controlled table extraction within a single framework. It features 30 diverse chart types and over 8k+ instances rendered via multiple libraries, providing rich image–table–code tuples for robust evaluation.
3 The ChartAnchor Benchmark
Figure 3: Illustration of the Data Collection Pipeline.
@@ 3 The ChartAnchor Benchmark
Table 1: Dataset statistics.
@@
Table 2: Chart counts across types.
ChartAnchor evaluates chart grounding via chart-to-code generation and controlled table reconstruction. We construct a diverse corpus of chart–table–code triplets from existing and real-world sources (distribution in Fig. 2). Our pipeline integrates parametric generation with augmentation, followed by a rigorous hybrid filtering process to ensure high data quality and semantic fidelity.
3.1 Task Definitions
As illustrated in Fig. 2, we define two core tasks grounded in our chart–table–code triplets:
Chart-to-Code Generation. This task requires generating executable Python code from a chart image $I$. The model must produce a valid script that accurately reconstructs visual and structural elements (e.g., chart type, data, layout, and style). This assesses capabilities in visual abstraction and symbolic reasoning, demanding output that is both syntactically correct and semantically faithful to the input.
Controlled Chart-to-Table Reconstruction. This task recovers tabular data under constrained header supervision. Given an image $I$ and headers $\mathcal{H}=\{h_{1},\dots,h_{n}\}$, the model generates a table $\mathcal{T}=[r_{1},\dots,r_{m}]$ aligning numerical values to corresponding headers. The inclusion of headers removes label ambiguity, enabling a more targeted and reliable evaluation of data fidelity.
3.2 Data Collection
This section presents the data collection pipeline for ChartAnchor, which comprises 8,068 samples: 1,535 (19%) augmented from existing datasets and 6,533 (81%) curated from real-world sources. Detailed procedures are provided in Appendix A.
3.2.1 Data Source
  • Existing datasets. We leverage chart–table pairs from PlotQA (Methani et al., 2020), DVQA (Kafle et al., 2018), FigureQA (Kahou et al., 2017), and Vistext (Tang et al., 2023). Given the absence of original code, we synthesize plotting scripts from metadata, applying systematic augmentations to visual and semantic attributes to ensure diversity and mitigate leakage.
3.2.2 Code Generation and Augmentation for Existing Datasets.
3.2.3 Filtering.
Our filtering process consists the following two stages, more details are in Appendix A.3.
Automated Filtering. To ensure data fidelity, we implement a four-stage pipeline: (a) Completeness, pruning incomplete data triples; (b) Structural Filtering, excluding unstructured formats such as rasterized tables or maps; (c) Deduplication, removing redundancies via exact matching of compact statistical signatures; and (d) Executability, discarding scripts that trigger runtime errors or rendering failures.
Human Filtering. We enforced rigorous quality control through a double-blind review process involving experts proficient in deep learning and data visualization. Samples were grouped by chart type and filtered within each group to ensure consistency across families. Independent reviewers evaluated the triples based on a standardized rubric for semantic accuracy, visual clarity, and stylistic diversity, with discrepancies resolved by a third adjudicator to guarantee high annotation reliability.
(a) Code Length Distribution.
(b) Image Entropy Distribution.
(c) Brightness vs. Contrast.
(d) Visual / Color Complexity.
Figure 4: Chart-level visual statistics.
3.3 Dataset Analysis
Dataset Types Test Set Input Format Output Format Multi-task Eval Visual Open Source Plot Library Data Eval Full Ground
Chart Benchmarks
ChartQA(Masry et al., 2022) 3 10k I+NL NL - -
PlotQA(Methani et al., 2020) 3 34k I+NL NL - -
Chart-to-text(Kantharaj et al., 2022b) 6 44k I+NL NL - -
OpenCQA(Kantharaj et al., 2022a) 5 1.2k I+NL NL - -
ChartSumm(Rahman et al., 2023) 3 84k I+NL NL - -
Charxiv(Wang et al., 2024) 18 1.32k I+NL NL - -
MMC(Liu et al., 2023) 6 2k I+NL NL - -
ChartX(Xia et al., 2024) 18 6k I+NL NL - -
ChartBench(Xu et al., 2023) 9 2.1k I+NL NL - -
Code Generation Benchmarks
HumanEval(Chen et al., 2021) - 164 Code Code - - - -
MBPP(Austin et al., 2021) - 500 NL+Code Code - - - -
MMCode(Li et al., 2024b) - 263 I+NL Code - - - -
MatPLotBench(Yang et al., 2024c) 13 100 NL Code Mixture
PLot2Code(Wu et al., 2024a) 15 132 I+NL Code Singal
ChartMimic(Yang et al., 2024b) 22 4.8k I+NL Code Singal
Design2Code(Si et al., 2024) - 484 I+NL Code Singal
ChartAnchor 30 8.1k I+NL Code+NL Mixture
Table 3: Comparison of benchmarks for chart understanding and code generation. “I” = image, “NL” = natural language. “Visual” assesses unannotated chart reasoning. “Plot Library” denotes the use of multiple plotting libraries in the chart-to-code task, while “Data Eval” indicates whether data fidelity is evaluated within the chart-to-code task.
@@ 3.3 Dataset Analysis
As detailed in Tab. 2, ChartAnchor encompasses 30 visualization families. Beyond canonical 2D plots, it covers diverse categories including advanced 3D charts, hierarchical diagrams, polar variants, and financial charts. The distribution mimics realistic usage patterns, balancing frequent types (e.g., bar, line) with long-tail complex charts. By integrating curated benchmarks with real-world code, ChartAnchor incorporates diverse plotting libraries (e.g., Matplotlib, Plotly) and styling idioms, establishing a robust benchmark for MLLM chart grounding. Detailed statistics and visual examples are provided in Appendices B and G.
3.4 Comparison with Other Benchmarks
Table 3 presents a comprehensive comparison between our proposed benchmark and a wide range of existing datasets related to chart understanding and code generation.
Early benchmarks (e.g., ChartQA, PlotQA, Chart-to-Text) focus on NL tasks like QA and summarization, without structured protocols to evaluate bidirectional alignment between chart visuals and their underlying data or specifications. More recent multi-task datasets such as ChartArxiv and ChartBench broaden the scope, yet still omit code-based grounding, limiting their ability to assess executable understanding of visual semantics.
Code generation benchmarks like HumanEval and MBPP target text-to-code synthesis and lack visual input, making them unsuitable for evaluating multimodal alignment. Conversely, multimodal datasets like Plot2Code and ChartMimic introduce image-to-code tasks but fail to explicitly assess numerical fidelity or data reconstruction.
@@ 4 Evaluation Metrics
1) Textual Components Match. We extract textual components (titles, axis labels, legends, annotations) and compare against reference charts to verify semantic and positional consistency.
2) Color Fidelity. We quantify perceptual color differences using the CIEDE2000 (Sharma et al., 2005) metric in the CIE Lab color space, which aligns with human vision sensitivity. This evaluation covers both static elements and dynamic color mappings. For multi-color comparisons, we apply the Hungarian algorithm to optimally pair generated and reference colors, minimizing the total perceptual deviation ($\Delta E_{00}$) across matched pairs. This ensures semantic alignment of color associations while preserving numerical fidelity.
@@
Text Color Type Layout
Legend Title Axis Label Annos Avg.
Proprietary Multimodal Large Language Models
GPT4o (OpenAI, 2024) 91.88 63.53 72.83 67.20 76.36 69.98 34.64 70.06 80.73 35.85 74.25 58.25
GPT-5 (OpenAI, 2025) 91.93 66.11 86.18 83.63 87.17 80.77 51.21 82.21 92.3 46.55 84.52 72.93
Claude-3-7-Sonnet (Anthropic, 2025a) 78.60 54.97 69.11 67.19 70.57 65.46 40.79 59.89 76.53 30.30 65.82 54.59
Claude-4-5-Sonnet (Anthropic, 2025a) 83.89 60.46 79.15 74.86 80.65 73.78 49.44 65.9 85.05 36.7 78.93 64.97
Gemini-3-Pro (Comanici et al., 2025) 83.86 60.31 79.84 76.97 83.25 75.09 48.32 71.98 87.38 41.58 75.33 66.61
Open-Source Multimodal Large Language Models
InternVL3-2B (Zhu et al., 2025) 28.97 18.75 17.41 17.63 20.88 18.67 8.98 16.01 21.55 13.56 17.93 15.75
Qwen2.5-VL-3B-Instruct (Yang et al., 2024a) 48.07 28.24 31.80 26.86 39.44 31.58 12.29 27.01 40.19 16.71 38.63 25.55
Gemma-3-4B-it (Team et al., 2025c) 66.20 42.29 38.72 36.41 57.50 43.73 18.74 33.53 60.46 24.21 47.48 36.13
DeepSeek-VL2-27B-A4B (Wu et al., 2024b) 7.98 5.93 5.53 6.38 5.41 5.81 2.7 6.08 5.54 3.25 4.51 4.65
DeepSeek-VL-7B (Lu et al., 2024) 32.05 17.59 20.41 17.62 23.19 19.70 9.56 17.37 24.83 7.74 20.35 15.84
LLaVA-v1.6-Mistral-7B (Li et al., 2024a) 15.32 8.74 8.37 4.96 4.91 6.75 6.03 7.79 5.05 4.76 4.39 6.07
Qwen2.5-VL-7B-Instruct (Yang et al., 2024a) 67.59 41.75 41.46 43.60 61.90 47.18 18.45 41.83 64.30 27.85 51.85 39.92
MiMo-VL-7B-SFT (Team et al., 2025b) 35.56 24.95 28.95 28.52 24.6 26.76 15.06 26.91 25.64 18.53 26.91 23.30
MiMo-VL-7B-RL (Team et al., 2025b) 41.51 30.19 34.79 34.01 30.26 32.31 18.76 32.77 31.96 22.92 29.9 28.10
MiniCPM-V-2.6-8B (Yao et al., 2024) 26.87 16.76 16.88 14.21 16.18 16.01 7.91 13.77 16.92 12.95 14.35 13.51
Qwen3-VL-8B-Instruct (Bai et al., 2025) 62.16 43.11 48.8 48.82 52.92 48.41 30.97 48.45 55.56 27.27 54.96 44.27
InternVL3-9B (Zhu et al., 2025) 69.19 44.75 47.48 45.06 59.57 49.20 22.06 40.72 66.20 29.20 55.14 41.47
GLM-4V-9B (GLM et al., 2024b) 46.04 26.55 27.86 24.55 43.78 30.68 12.02 23.32 45.69 15.51 36.66 25.44
CogVLM2-Llama3-Chat-19B (Hong et al., 2024) 7.82 7.12 6.03 4.79 1.25 4.80 3.77 5.61 1.21 3.52 6.15 3.78
InternVL3-14B (Zhu et al., 2025) 80.60 52.80 55.69 54.90 75.12 59.63 26.13 53.08 80.63 30.59 64.7 50.01
Qwen2.5-VL-32B-Instruct (Yang et al., 2024a) 67.59 42.71 48.69 48.91 73.55 53.47 23.56 48.29 82.08 21.14 66.58 45.70
Qwen3-VL-32B-Instruct (Bai et al., 2025) 73.56 51.15 46.67 50.64 66.39 53.71 38.51 58.46 70.14 33.86 63.87 53.09
GLM-4.5V-106B (Team et al., 2025a) 66.77 49.18 61.44 58.12 55.49 56.06 32.55 55.54 57.97 34.03 52.58 48.12
Qwen3-VL-235B-A22B-Instruct (Bai et al., 2025) 79.62 57.23 64.54 62.65 73.67 64.52 37.2 62.88 76.81 38.07 68.62 58.02
Table 4: Comprehensive performance comparison of proprietary and open-source MLLMs on the Chart-to-Code generation task. For data fidelity, we report IoU score under a slight tolerance stage. As our collected matplotlib images do not include annotations by design, they are excluded from the annotation metric evaluation.
To evaluate whether models can faithfully recover the underlying chart data, we introduce data-level fidelity metricfor both chart-to-code and controlled chart-to-table tasks.
@@
In the chart-to-code, we first parse the generated figure objects and dispatch them to type-specific extractors $\text{Extract}_{T}$, where each $T\in\mathcal{T}$ corresponds to a supported chart type. Each extractor retrieves the relevant data fields—such as $(x,y)$ for line or $(x,\text{low},\text{high},\text{open},\text{close})$ for candlestick charts—and formats each data point as a normalized tuple $\tau_{i}=(n,x_{i},y_{i},\ldots)$, where $n$ is the trace name and the remaining elements are type-specific field values. The final structured form is a set $\mathcal{L}_{T}=\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$. In the controlled chart-to-table task, each table row is similarly treated as a tuple $\tau_{i}=(t_{1},t_{2},\ldots,t_{m})$, resulting in a comparable structured set $\mathcal{L}_{T}$, allowing both tasks to be evaluated under a unified tuple-based framework.